Three pipes can fill a pool in eight hours.A drain can empty the fill pool in twelve hours.How many hours will it take to fill a quarter of the fool if the three pipes and drain open?
Answer by stanbon(75887)

Three pipes can fill a pool in eight hours.A drain can empty the fill pool in twelve hours.How many hours will it take to fill a quarter of the fool if the three pipes and drain open?
-----
3-pipe rate: 1/8 job/hr
Drain rate: 1/1/12 job/hr
Together rate: 1/x job/hr
-----
rate - rate = together rate
1/8 - 1/12 = 1/x
12x - 8x = 8*12
4x = 8*12
x = 24 hrs (time to fill the pool together)
======================
Cheers,
Stan H.
================